OpenStudy (anonymous):

you roll 2 dice. what is the probability that the sum of the dice is even OR the sum of the dice is less than 6? Probability is tricky for me!

OpenStudy (anonymous):

Trick For you...

OpenStudy (anonymous):

I think I figured it out: is it 2/3?

OpenStudy (anonymous):

the sum of the dice is even ={(1,1),(1,3),(1,5),(3,1),(3,3)(3,5),(5,1)(5,3),(5,5),(2,2),(2,4),(2,6),((4,2),(4,4),(4,6),(6,2),(6,4),(6,6)} OR the sum of the dice is less than 6 = {(1,1),(1,2),(1,3),(1,4),((2,1),(2,2),(2,3),(3,2),(4,1)} ,,,, p (the sum of the dice is even )= 18/36 = 1/2 , ,,,,,,,,, p(the sum of the dice is less than 6) = 9/36=1/4 ,,,,,,,,, P(A and B) = 3/36=1/12 p(A or B) = p(a)+P(B) -P(a and B) = 1/2 +1/4 - 1/12 = 2/3